OpenStudy (anonymous):

Which expression is the simplified form of ? Which expression is the simplified form of ? @Mathematics

OpenStudy (anonymous):

\[\sqrt{k^23\sqrt{?}}\]

OpenStudy (anonymous):

i dont think i understand the expression...

OpenStudy (anonymous):

Its the square root sign then under it is k^23

OpenStudy (anonymous):

u mean this?:\[\sqrt{k^{23}}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you could write the root like this:\[\sqrt{k^{23}} = k^{\frac{23}{2}} = (k^{11} \times k^{\frac{1}{2}}) = k^{11}\sqrt{k}\]

OpenStudy (anonymous):

Which expression is the simplified form of \[\sqrt{63w^36}\] ?

OpenStudy (anonymous):

the 6 on the end means 36

OpenStudy (anonymous):

\[\sqrt{63w^{36}} = \sqrt{63}w^{\frac{36}{2}} = w^{18}\sqrt{63}\]looking at 63^(1/2)\[\sqrt{63} = \sqrt{21 \times 3} = \sqrt{7 \times 3 \times 3} = 3\sqrt{7}\]then you get\[3w^{18}\sqrt{7}\]

OpenStudy (anonymous):

Ok i have a couple more can you still help me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\sqrt{5+3 \sqrt {10+2\sqrt{5}}}\]What is the exact value of the expression

OpenStudy (anonymous):

\[3\sqrt{10+\sqrt{20}} = \sqrt{90 + 9\sqrt{20})} = \sqrt{90 + \sqrt{1620}}\]hmmm.... i think u need a calculator for this one -.-

OpenStudy (anonymous):

well i got like 4.051256372 from the calculator >.>

OpenStudy (anonymous):

\[\sqrt{72 - \sqrt{8+\sqrt{128}}}\]What is the exact value of the expression

OpenStudy (anonymous):

8.2222402002 i think u can use a calculator for solving these, just watch out for the order of operations